OpenStudy (anonymous):

Solve the equation for h: S= lw + wh + hl Would I solve for both h's or just one?

OpenStudy (anonymous):

well h=h so you will only need to solve for one

OpenStudy (ivettef365):

you need to isolate the h, so first step is to move the lw, to the other side

OpenStudy (ivettef365):

so you have S - lw = wh + hl

OpenStudy (ivettef365):

now you can write the problem as S - lw = h(w+l)

OpenStudy (anonymous):

K so I can either solve for the h in wh or hl..?

OpenStudy (ivettef365):

now just divide by w + l and you have S - lw ------ = h w+l

OpenStudy (anonymous):

Ohhh I didn't look at it that way... Thanks!

OpenStudy (ivettef365):

yw :)